Kind, Inc., an exempt organization, leases land, a building, and factory equipment to Shirts, Inc. Shirts is a taxable entity that manufactures shirts for distribution through its factory outlet stores. The rent income and the related expenses for Kind are as follows:

Rent Income Rent Expenses
Land and building $100,000 $40,000
Factory equipment 125,000 25,000

a. The amount of Kind's unrelated business income is $.

b. Assume instead that Kind's rent income and expenses are recorded as follows.

Rent Income Rent Expenses
Land and building $100,000 $20,000
Factory equipment 125,000 50,000

The amount of Kind's unrelated business income is $.

Please answer a and b with explanation. Not sure if there would be a difference with the rent expenses. [Taxation 2019 Question] Thank You!
Answer #1

(a)  

Rent Income Rent Expenses
Land and building $100,000 $40,000
Factory equipment 125,000 25,000

Land and building rent net income= 100000 - 40000 = $ 60000

Factory Equipment rent net income = 125000-25000 = $ 100000

Unrelated business income = 60000 + 100000 = $ 160000

As kind is exempt organization so rent is not an business activity done in normal course all these will be unrelated business income = $ 160000

Rent Income Rent Expenses
Land and building $100,000 $20,000
Factory equipment 125,000 50,000

(b)

Land and building rent net income= 100000 - 20000 = $ 80000

Factory Equipment rent net income = 125000-50000 = $ 75000

Unrelated business income = 80000 + 75000 = $ 155000

As kind is exempt organization so rent is not an business activity done in normal course all these will be unrelated business income = $ 155000